1. If I have 5.6 liters of gas in a piston at a pressure of 1.5 atm and compress the gas until its volume is 4.8 L, what will the new pressure inside the piston be?

2. I have added 15 L of air to a balloon at sea level (1.0 atm). If I take the balloon with me to Denver, where the air pressure is 0.85 atm, what will the new volume of the balloon be?

3. I’ve got a car with an internal volume of 12,000 L. If I drive my car into the river and it implodes, what will be the volume of the gas when the pressure goes from 1.0 atm to 1.4 atm?

  1. Some students' believe teachers are full of hot air. If I inhale 2.2 L of air at 18 0C, if it heats to a temperature of 38 0C in my lungs, what volume will it occupy?
Answer #1

1. From Boyle's law,

P1 x V1 = P2 x V2

Given that, P1 = 1.5 atm , V1 = 5.6 L , P2 = ? , V2 = 4.8 L

=> 1.5 atm x 5.6 L = P2 x 4.8 L

=> P2 = 1.75 atm

2. From Boyle's law,

P1 x V1 = P2 x V2

Given that, P1 = 1.0 atm , V1 = 15 L , P2 = 0.85 atm , V2 = ?

=> 1.0 atm x 15 L = 0.85 atm x V2

=> V2 = 17.65 L

3. From Boyle's law,

P1 x V1 = P2 x V2

Given that, P1 = 1.0 atm , V1 = 12000 L , P2 = 1.4 atm , V2 = ?

=> 1.0 atm x 12000 L = 1.4 atm x V2

=> V2 = 8571.4 L

4. From Charles law,

V1 / V2 = T1 / T2

Given that, V1= 2.2 L , T1 = 18 C = 18+273 = 291 K , V2 = ? , T2 = 38 C = 38 + 273 = 311 K

=> 2.2 L / V2 = 291K / 311K

=> V2 = 2.35 L